40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 25 02:04:19 2014, Jenne, Update, LSC, PRFPMI tonight - need some PRCL and MICH tuning at high arm powers HighPower_7pt3sec_Powers.pngHighPower_7pt3sec_ErrAndCtrls.pngHighPower_7pt3sec_AuxErrs.pngCARMDARMonALScommdiff_StayingAbove5_upTo175_24Nov2014.png
    Reply  Wed Nov 26 05:15:48 2014, Jenne, Update, LSC, PRFPMI tonight PRMI 100Hz osc? PRFPMI_currentErrSigs_25Nov2014.pdf
Message ID: 10734     Entry time: Tue Nov 25 02:04:19 2014     Reply to this: 10736
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRFPMI tonight - need some PRCL and MICH tuning at high arm powers 

Take-away for the night:  We need to do some more fine-tuning of the PRCL and MICH loops when we have arm resonance.

Koji sat with me for the first part of the night, and we looked back at the data from last week (elog 10727), as well as some fresh data from tonight.  Looking at the spectra, we noticed that last week, and early in the evening today, I had a fairly broad peak centered around ~51Hz.  We are not at all sure where this is coming from.  The PRMI was locked on REFL 33 I&Q, and CARM and DARM were both on ALS comm and diff.  This peak would repeat-ably come and go when I changed the CARM offset.  At high arm powers (above a few tens? I don't know where exactly), the peak would show up.  Move off resonance, and the peak goes away.  However, later in the night, after an IFO realignment, I wasn't able to reproduce this effect.  So.  We aren't sure where it comes from, but it is visible only in the CARM spectra, so there's some definite feedback funny business going on. 

Anyhow, after that, since I couldn't reproduce it, I went on to trying to hold the PRMI at high arm powers, but wasn't so successful.  I would reduce the CARM offset, and instead of a 50Hz peak, I would get broadband noise in the PRMI error signals, that would eventually also couple in to the CARM (but not DARM) error signal, and I would lose PRMI lock. I measured the PRCL and MICH transfer functions while the arms were at some few units of power, and found that while MICH was fine, PRCL was losing too much phase at 100Hz, so I took away the FM3 boost.  This helped, but not enough. I had 1's in the triggering matrix for TRX and TRY to both PRCL and MICH, so that even if POP22 went low, if the arms were still locked then the PRMI wouldn't lose lock unnecessarily, but I was still having trouble.  In an effort to get around this, I transitioned PRMI over to REFL 165 I&Q. 

While the arms were held around powers of 2ish, I readjusted the REFL 165 demod phase.  I found it set to 150 deg, but 75 deg is better for PRMI locking with the arms.  For either acquiring or transitioning from REFL33, I would use REFL165I * -1.5 for PRCL, and REFL 165Q * 0.75 for MICH.  (Actually, I was using -2 for REFL165I->PRCL, and +0.9 for REFL165Q->MICH, but I had to lower the servo gains, so doing some a posteriori math gives me -1.5 and +0.75 for what my matrix elements should have been, if I wanted to leave my servo gains at 2.4 for MICH and -0.02 for PRCL.) I don't always acquire on REFL165, and if it's taking a while I'll go back to putting 1's in the REFL33 I&Q matrix elements and then make the transition. 

With PRMI on REFL 165 I&Q, I no longer had any trouble keeping the PRMI locked at arbitrarily high arm powers.  I was still using 1*POP22I + 1*TRX + 1*TRY for triggering PRCL and MICH.  My thresholds were 50 up, 0.1 down.  The idea is that even if POP goes low (which we've seen about halfway up the CARM resonance), if we're getting some power recycling and the arms are above 1ish, then that means that the PRMI is still locked and we shouldn't un-trigger anything.  I didn't try switching over to POP110 for triggering, because POP22 was working fine.

Earlier in the night, Koji and I had seen brief linear regions in POX and POY, as well as some of the REFL signals when we passed quickly through the CARM resonance.  I don't have plots of these, but they should be easy to reproduce tomorrow night.  Koji tried a few times to blend in some POY to the CARM error signal, but we were not ever successful with that.  But, since we can see the PDH-y looking regions, there may be some hope, especially if Q tells us about his super secret new CESAR plan. 

Okay, I'm clearly too tired to be writing, but here are some plots.  The message from these is that the PRMI loops are causing us to fluctuate wildly in arm transmission power.  We should fix this, since it won't go away by getting off of ALS.  The plots are from a time when I had the PRMI locked on REFL165, and CARM and DARM were still on ALS comm and diff.  All 3 of these colored plots have the same x-axis.   They should really be one giant stacked plot.

HighPower_7pt3sec_Powers.png

HighPower_7pt3sec_ErrAndCtrls.png

HighPower_7pt3sec_AuxErrs.png

Also, bonus plot of a time when the arm powers went almost to 200:

CARMDARMonALScommdiff_StayingAbove5_upTo175_24Nov2014.png

ELOG V3.1.3-